Using these data from the comparative balance sheet of Rollaird Company, perform horizontal analysis.

December 31, 2020

December 31, 2019

Accounts receivable

$545,300

$410,000

Inventory

806,200

580,000

Total assets

3,052,660

2,587,000


(Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

Increase or (Decrease)

December 31, 2020

December 31, 2019

Amount

Percentage

Accounts receivable

$545,300

$410,000

$

%

Inventory

806,200

580,000

%

Total assets

3,052,660

2,587,000

%
Answer #1

Answer

Rollaird company

Comparative Balance sheet

Increase or (Decrease)
Particulars 2020 2019 Amount Percentage
Account receivables 545,300 410,000 135,300 33%
Inventory 806,200 580,000 226,200 39%
Total assets 3,052,660 2,587,000 465,660 18%

Working note :

Change in Amount

Account receivables = 545,300 - 410,000 = $135,300

Inventory = 806,200 - 580,000 = $226,200

Total Assets = 3,052,660 - 2,587,000 = $465,660

Change in percentage

Account receivables = (135,300 / 410,000) × 100 = 33%

Inventory = (226,200 / 580,000) × 100 = 39%

Total Assets = (465,660 / 2,587,000) × 100 = 18%

Account receivables changed 33% , total assets change 18% and inventory change 39%.

If YOU are like the answer please upvote.